# A detailed study on ergodic to non-ergodic phase transition in the dissipative anisotropic Dicke model

### v1

@@ III.1 Finite size scaling of the Liouvillian gap
where $E_{n}$ are the eigenvalues of the liouvillian and $E_{0}=0+0i$, is the corresponding minimum eigenvalue. In Fig. 2 we study the Liouvillian gap as a function of the total dimension of the Liouvillian space for $g_{1}=1.25$, and $g_{2}=0.1-1.5$, ranging from the non-ergodic/regular phase to the ergodic/quantum chaotic phase. We notice that in the ergodic phase, $\Delta\propto N^{-z}_{L}$ with $10<z<$ 1, whereas in the non-ergodic phase it is independent of the total Liouvillian space dimension. We also plot the same data for $\Delta$ as a function of the atom number $N$, (for the fixed $n_{\text{max}}$) and see a similar behavior (scaling) as the previous case. We note that our scaling result is similar to one of the cases of bulk dissipation highlighted recently in the work by Znidaric. Žnidarič (2015). The reason for this type of scaling is given below.
In the ergodic phase, the system dynamics occurs in the full Hilbert space, so by increasing the number of atoms, the time to cover the full space also increases. The negative slope in the plots indicates that the relaxation time increases with the atom number. On the other hand, in the case of the non-ergodic phase, due to the presence of symmetry and conserved quantities, the dynamics takes part in a relatively smaller subspace. Hence for a higher number of atoms, no such changes in the $\Delta$ is noticed here. We consider only the spontaneous emission process in the dissipative ADM and hence it shows different behavior from the result of the previous work considering the dephasing Mori (2024).
However, we confine ourselves to the finite number of atoms and observe that in this limit, the system has a finite $\Delta$. Our future focus will be on the behavior of ($\Delta$) in the thermodynamic limit.
III.2 Eigenstate properties
The eigenvalue properties of the dissipative Dicke model are well studied Prasad et al. (2022). We also provide a brief discussion of the eigenvalue properties in the Appendix. A recent study exists in the literature that includes the anisotropic case of this particular model Villaseñor et al. (2024). However, a complete study requires both the eigenvalue and eigenvector properties to understand the phase diagram of this model. In this section, we discuss the eigenstate property and check its consistency with the eigenvalue property.
III.2.1 Average participation ratio
It is well known that analyzing the eigenstates of non-Hermitian systems can predict the emergence of chaos, e.g., many-body coupled Hatano-Nelson chain Suthar et al. (2022). Motivated by the above approach, we study the eigenstate properties of the Lindbladian super-operator. In particular, we focus on the average participation ratio (PR), $P_{\text{avg}}$. Similar to the non-Hermitian Hamiltonian, the right and left eigenstates are not the same. Therefore, we perform a bi-orthogonalization to make an orthogonal basis that can span the entire Liouvillian space, which is given below in detail. We also note that the eigenstates with negative eigenvalues have a finite lifetime Saha and Bhattacharyya (2024). Hence, our results can only be envisaged as the chaotic properties of the non-equilibrium states which may not have an equilibrium counterpart.
@@ III.2.1 Average participation ratio
In the case of a Liouville space, we perform the LU decomposition for bi-orthogonalization. Let $L$ and $R$ be the corresponding left and right eigenvectors of $\mathcal{L}$ and $L^{\dagger}R=M$. The matrix $M$ can be represented in its LU-decomposed form: $M=M_{\text{L}}M_{\text{U}}$, where $M_{\text{L}}(M_{\text{U}})$ is the lower(upper) triangular matrix. Hence,
$L^{\dagger}R=M_{\text{L}}M_{\text{U}}$
or
$M_{\text{L}}^{-1}L^{\dagger}RM_{\text{U}}^{-1}=1.$ (7)
@@
Following the above approach, one can redefine the left and right eigenvectors as $L\rightarrow L\left(M_{\text{L}}^{-1}\right)^{\dagger}$ and $R\rightarrow RM_{\text{U}}^{-1}$ such that $L^{\dagger}R=1$ to make the Liouvillian space bi-orthogonal.
For non-Hermitian systems, PR can be defined in two ways: first, using the $n$th left or right eigenstate, and the second is defined using the biorthogonal basis from both left and right eigenstates. The definition of PR is given by:
$P_{n}=\frac{\left(\sum_{i}|\psi_{i}^{n}|\right)^{2}}{\sum_{i}|\psi_{i}^{n}|^{2}}$ (8)
$P_{nB}=\frac{\left(\sum_{i}|\tilde{\psi}_{i}^{n}|\right)^{2}}{\sum_{i}|\tilde{\psi}_{i}^{n}|^{2}}.$ (9)
The subscript $nB$ represents biorthogonal PR, $\psi_{i}^{n}=(\langle R_{n}|b_{i}\rangle^{*})\langle R_{n}|b_{i}\rangle$ and $\tilde{\psi}_{i}^{n}=(\langle L_{n}|b_{i}\rangle^{*})\langle R_{n}|b_{i}\rangle$ with $R_{n}$ and $L_{n}$ are the $n^{\text{th}}$ right and left eigenstates respectively, and $|b_{i}\rangle$ are Fock space chosen as computational basis. The average PR is defined as: $P_{\text{avg}}=\sum_{n}P_{\text{n}}/N_{\text{L}}$ and $P^{B}_{\text{avg}}=\sum_{n}P_{\text{nB}}/N_{\text{L}}$. In Fig. 3 we show the value of $P_{\text{avg}}$ and $P_{\text{avg}}^{B}$ as a function of the coupling strengths. In the dark portion, the average participation ratio is relatively low which indicates the non-ergodic phase, compared to the light color regime where the average participation is quite high and represents the ergodic phase. The phase diagrams are similar to Fig. 9 in the Appendix, where we show the transition from the non-ergodic to the ergodic phase with the aid of level spacing ratio. Hence we can say that the eigenvector properties of $\mathcal{L}$ in our system are consistent with the eigenvalue properties. In Fig. 4 we highlight the data along the diagonal direction $g_{1}=g_{2}=g$ which indicates the isotropic Dicke model. Here the vertical dot-dashed line represents the critical coupling strength of the phase transition at $g=g^{*}=0.7071$. The variation between the three different types of participation ratios in Fig. 3 is interesting and worthy of a closer investigation.
@@
Figure 5: (a, b, e, f, i, j, m, n) Average boson number, $N_{\text{av}}(t)$, and (c, d, g, h, k, l, o, p) mutual information between spins and bosons, $I_{\text{spin-boson}}$, as a function of time $t_{n}=2^{n}T$ for the dissipative ADM under the Thue-Morse quasiperiodic drive. The initial states have low energies, so that $\langle E_{\text{in}}\rangle=0.25$. Panels (a, c, e, g, i, k, m, o) represent the data for a fixed driving frequency $\omega_{\text{d}}=50$, $g_{1}=1.25$, and various values of $g_{2}$. Panels (b, d, f, h, j, l, n, p): $g_{1}=1.25$, $g_{2}=0.7$ and various values $\omega_{\text{d}}$. In all panels, the driving amplitude is $\Omega=1.0$, $N=6,\ n_{\text{max}}=20,\ \omega_{0}=\omega=1$. For (a-d) $\kappa=0.0$, (e-h) $\kappa=0.05$, (i-l) $\kappa=0.6$, (m-p) $\kappa=1.0$. The time t is in units of $\omega_{0}^{-1}$.
IV Analysis of the driven dissipative ADM
In our earlier study of driven closed ADM Das et al. (2023b), we show that under the external periodic drive, the system reaches a plateau that escapes Floquet heating. In contrast, for the quasiperiodic drive, the system features a prethermal plateau and finally reaches an infinite temperature state. Such analysis can also be extended for the dissipative case. From the analysis of the previous sections, we note that the non-ergodic phase becomes more stable in the presence of bosonic dissipation. Whereas, our previous evidence shows that the quasiperiodic drive pushes the system to the ergodic phase. Hence, initiating the competition between the quasiperiodic drive and the bosonic dissipation to stabilize the non-ergodic phase is the main focus of this section.
@@ IV Analysis of the driven dissipative ADM
$Driving\ protocol\ -$ We briefly discuss the quasi-periodic driving protocol. In particular, we consider the quasi-periodic protocol consisting of the Thue-Morse sequence. Here, we define two Hamiltonian $H_{\pm}=\omega a^{\dagger}a+\omega_{0}J_{z}+\frac{(g_{1}\pm\Omega)}{\sqrt{N}}(aJ_{+}+a^{\dagger}J_{-})+\frac{(g_{2}\pm\Omega)}{\sqrt{N}}(aJ_{-}+a^{\dagger}J_{+})$, where $\Omega$ is the driving amplitude and the corresponding Liouvillian using Eq. 4 are $\mathcal{L}_{\pm}$. The Thue-Morse sequence Thue (1906); Nandy et al. (2017); Mukherjee et al. (2020); Zhao et al. (2021, 2022) can be constructed by starting time evolution operators $\mathcal{U}_{\pm}=\exp(\mathcal{L}_{\pm}T)$. We can build up the next sequence using $\mathcal{U}_{1}=\mathcal{U}_{-}\mathcal{U}_{+}$ and $\tilde{\mathcal{U}}_{1}=\mathcal{U}_{+}\mathcal{U}_{-}$, $\mathcal{U}_{2}=\tilde{\mathcal{U}}_{1}\mathcal{U}_{1}$ and $\tilde{\mathcal{U}}_{2}=\mathcal{U}_{1}\tilde{\mathcal{U}}_{1}$ and so on. The driving unit cells of time length $2^{n}T$ can be recursively constructed as: $\mathcal{U}_{n+1}=\tilde{\mathcal{U}}_{n}\mathcal{U}_{n},\ \tilde{\mathcal{U}}_{n+1}=\mathcal{U}_{n}\tilde{\mathcal{U}}_{n}$. The time evolution of the initial density matrix is given by $\rho_{n}|_{N_{\text{L}}\times 1}=\rho(t=2^{n}T)|_{N_{\text{L}}\times 1}=\mathcal{U}_{n}|_{N_{\text{L}}\times N_{\text{L}}}\times\rho_{\text{in}}|_{N_{\text{L}}\times 1}$, where $\rho_{\text{in}}|_{N_{\text{L}}\times 1}=|\psi_{\text{in}}\rangle|_{N_{\text{D}}\times 1}\otimes|\psi_{\text{in}}\rangle|_{N_{\text{D}}\times 1}$ is the initial density matrix and $|\psi_{\text{in}}\rangle$ is the initial state. To calculate the expectation value of an observable we reshape the density matrix as: $\rho(t)|_{N_{\text{L}}\times 1}\rightarrow\rho(t)|_{N_{\text{D}}\times N_{\text{D}}}$. Here we take the average over $20$ low energy eigenstates (starting from the ground state) of the decoupled Hamiltonian ($g_{1}=g_{2}=0.0$).
IV.1 Quantum dynamics
@@
To study quantum dynamics we consider the average boson number:
$N_{\text{av}}(t)=\text{Tr}\left[\rho(t)|_{N_{\text{D}}\times N_{\text{D}}}\times(a^{\dagger}a)|_{N_{\text{D}}\times N_{\text{D}}}\right]$ (10)
@@
where $S_{\text{spin-boson}}=-\text{Tr}\left[\rho(t)\ln(\rho(t))\right]$, $S_{\text{spin(boson)}}=-\text{Tr}\left[\rho_{\text{spins(boson)}}(t)\ln(\rho_{\text{spins(boson)}}(t))\right]$ are the corresponding entropies and $\rho_{\text{spin}}(t)|_{(N+1)\times(N+1)}=\text{Tr}_{\text{boson}}\left[\rho(t)|_{N_{\text{D}}\times N_{\text{D}}}\right]$, $\rho_{\text{boson}}(t)|_{(n_{\text{max}}+1)\times(n_{\text{max}}+1)}=\text{Tr}_{\text{spin}}\left[\rho(t)|_{N_{\text{D}}\times N_{\text{D}}}\right]$ are the corresponding density matrix or reduced density matrix of the spin(boson) obtained by tracing over the bosonic(spin) degrees of freedom.
In Fig. 5, we consider low-energy initial states and the Thue-Morse driving protocol. We show the dynamics of the average boson number [Fig. 5(a, b, e, f, i, j, m, n)] and the mutual information between spins and bosons [Fig. 5(c, d, g, h, k, l, o, p)] for a fixed intermediate value of the driving frequency $\omega_{\text{d}}$ and various values of the coupling parameter $g_{2}$ [Fig. 5(a, c, e, g, i, k, m, o)]. We also study the same quantities for a fixed $(g_{1},g_{2})$ pair associated with the chaotic undriven model and various values of $\omega_{\text{d}}$ [Fig. 5(b, d, f, h, j, l, n, p)]. For the close system ($\kappa=0$) von Neumann entanglement entropy is a good quantifier as the spin-boson density matrix, $\rho(t)$ forms a pure state. On the other hand, when we switch on $\kappa$, $\rho(t)$ doesn’t remain a pure state, it becomes a mixed state. As mutual information is a good quantifier for analyzing a mixed state, we pick this observable for our study. Fig. 5(a-d) correspond to $\kappa=0$ and it shows a result similar to the closed quasiperiodically driven system. There are three characteristic features: an initial increasing part followed by an intermediate prethermal plateau and a finial infinite temperature state. On the other hand, all other panels with $\kappa\neq 0$ exhibit an initial increasing part, and then the system remains in a prethermal plateau which is not followed by the final infinite-temperature state, and this contrasts with the results in the closed ADM ($\kappa=0$ case), (see the Fig. 5(a-d) here, or the Fig. 4 in our earlier work Das et al. (2023b)). The reason behind the nature of the $\kappa\neq 0$ case is as follows: in a closed-driven system, the system absorbs energy from the drive and tends to reach an infinite-temperature state over time. At very high driving frequencies, this can lead to the emergence of a prethermal phase Das et al. (2023b). On the other hand, the bosonic dissipation leads the boson to the ground states, which lowers the entropy of the system.
For a better understanding of the effect of dissipation, we also provide the dynamical equation of the spins at the strong dissipation limit where the bosonic degrees of freedom can be adiabatically eliminated. For $\kappa>>g_{1},\ g_{2}$, the dynamical equation of the spins is given by Kessler et al. (2012); Link and Strunz (2020). ,
$\dot{\rho_{s}}=-i[\omega J_{\text{z}},\rho_{s}]+\kappa_{\rm eff}[2J_{-}\rho J_{+}-\{J_{+}J_{-},\rho_{s}\}].$ (12)
Here $\rho_{s}$ is the spin density matrix. The above dynamical equation represents the Dicke model of resonance fluorescence. We note that the quasi-periodic drive has no effect (in $\kappa>>g_{1},\ g_{2}$ limit) and the steady state will be the ground state.
Another interesting feature in the high $\kappa$ regime is the linear dependency of $\Delta$ on atom number $N$, which is well known for the Dicke superradiance burst and shows a sharp contrast from the finite values of $\kappa$ Gross and Haroche (1982).
V Conclusion
[…]
We also extend the analysis to the driven case where a quasiperiodic Thue-Morse sequence is applied to the system and study the quantum dynamics with the aid of average boson number and mutual information between spins and bosons in the system. Here we see the observables are initially increasing with the sequential time followed by a stabilized prethermal plateau. The system is not going to the final infinite temperature state which contrasts the result for the closed anisotropic quasiperiodically driven system and not reported previously.
Effect of the dissipation and dephasing in the superradiance phase transition was already studied in previous works by Kirton et al Kirton and Keeling (2017). Our work provides an addition to the field of the dissipative Dicke model, where the non-ergodic phase transition is thoroughly studied and analyzed for both the driven and undriven cases.
Further investigation considering the individual decay processes of the atoms is required to understand the other illuminating aspects of the dissipative Dicke model. The dimensionality of the Liouvillian is an issue for simulating large numbers of atoms, hence future research works demand advanced numerical techniques to overcome this difficulty Weimer et al. (2021).

### v2

@@ III.2.1 Finite size scaling of the Liouvillian gap
where $E_{n}$ are the eigenvalues of the liouvillian and $E_{0}=0+0i$, is the corresponding minimum eigenvalue. The inverse of the Liouvillian gap, $1/\Delta$, shows the notion of the relaxation time of the system. In Fig. 1(a), we plot $\ln\Delta$ versus $\ln N$ for fixed $g_{1}=1.0$ and $\kappa=0.1$, and for different values of $g_{2}$ ($0.1\leqslant g_{2}\leqslant 1.5$). From the linear fits in Fig. 1(a), we extract the scaling exponent $z$ defined by $\Delta\propto N^{z}$. The resulting dependence of $z$ on $g_{2}$ is shown in Fig. 1(b). We find that the slope of $z$ is different for the non-ergodic phase and the ergodic phase, respectively. The system exhibits a non-ergodic response when $g_{2}$ is small, while increasing $g_{2}$, we find a crossover into the ergodic phase. A more thorough analysis is required to reliably identify the phase boundary of ENET, in a manner similar to that used for characterizing a QPT.
A true phase transition would require taking the thermodynamic limit $N\to\infty$; however, the high dimensionality of the system makes this limit challenging to access. Hence, our result shows a phase crossover between the ergodic and non-ergodic phases rather than a phase transition.
We check the same for $g_{1}=1.0,\,1.25$ and $\kappa=0.1,\ 0.01$ in Fig. 1(b), and our result shows that for changing $\kappa$, the system shows qualitatively similar response. We also provide a contour plot of the exponent $z$ by varying $g_{1}$ and $g_{2}$ in Appendix B to provide other details on ENET.
Figure 1: (a) $\ln\Delta$ vs $\ln N$ plot for $g_{1}=1.0$, and $\kappa=0.1$. (b) The slope $z$ (extracted from the scaling of Liouvillian gap $\Delta$ as a function of atom number) as a function of $g_{2}$ for fixed $g_{1}$ value. We consider the dissipation strength, $\kappa=0.1,0.01$. Here, $n_{\text{max}}=26$, and the atom number NN is changing from 4104-10. We find that the scaling is different for the ergodic and non-ergodic phases. We also show that for changing $\kappa$, the qualitative features of ENET remain the same.
III.2.2 Average participation ratio using eigenvector properties
The eigenvalue properties of the Liouvillian for the dissipative Dicke model are well studied previously  Prasad et al. (2022). A recent study including the anisotropic case of this particular model also exists  Villaseñor et al. (2024). However, for completeness, a study requires both the eigenvalue and the eigenvector properties to understand the phase diagram of this model. We also note that for non-Hermitian systems, the correspondence between the eigenvalue and eigenvector is not as straightforward as in the Hermitian case. For example, if the Liouvillian of the dissipative system has a non-zero real eigenvalue, then the corresponding eigenvector is traceless, which is not true for the Hamiltonian of the Hermitian case. This observation motivates us to study the eigenvector properties of the dissipative ADM in order to gain deeper insight into ENET.
In general, Liouvillian is a non-Hermitian matrix with complex eigenvalues, $\mathcal{L}\rho_{i}=\lambda_{i}\rho_{i}$. If $\mathcal{L}$ is diagonalizable, then using the spectral decomposition theory, any operator $\hat{A}$ in the Liouville space can be written as $\hat{A}=\sum c_{i}\rho_{i}$. The value of $c_{i}$ can be calculated from the standard orthogonality relation for a non-Hermitian matrix, $Tr(\Theta_{j}^{\dagger}\rho_{i})=\delta_{ij}$. Here, $\Theta_{j}$ is the left eigenvector of $\mathcal{L}$, i.e., $\mathcal{L}^{\dagger}\Theta_{i}=\lambda_{i}\Theta_{i}$. The above definition implies that to calculate the participation ratio $PR\propto\frac{1}{\sum_{i=1}^{N}|c_{i}|^{4}}$ Das et al. (2023a), we must use a bi-orthogonal basis consisting of both left and right eigenvectors. We note that the results corresponding only to the left and only to the right are not quite correct and may provide wrong results.
@@ III.2.2 Average participation ratio using eigenvector properties
To calculate the participation ratio, in the Liouville space, we perform the LU decomposition for bi-orthogonalization. Let $L$ and $R$ be the corresponding left and right eigenvectors of $\mathcal{L}$, and we define $L^{\dagger}R=M$. The matrix $M$ can be represented in its LU-decomposed form: $M=M_{\text{L}}M_{\text{U}}$, where $M_{\text{L}}(M_{\text{U}})$ is the lower(upper) triangular matrix. Hence,
$L^{\dagger}R=M_{\text{L}}M_{\text{U}}$ (7)
@@
Following the above approach, one can redefine the left and right eigenvectors as $L\rightarrow L\left(M_{\text{L}}^{-1}\right)^{\dagger}$ and $R\rightarrow RM_{\text{U}}^{-1}$ such that $L^{\dagger}R=1$ to make the Liouvillian space bi-orthogonal. The definition of PR is given by:
$P^{\rm LR}_{n}=\frac{\left(\sum_{i}|\tilde{\psi}_{i}^{n}|\right)^{2}}{\sum_{i}|\tilde{\psi}_{i}^{n}|^{2}}.$ (8)
The superscript $\rm LR$ represents biorthogonal PR. $\tilde{\psi}_{i}^{n}=(\langle L_{n}|b_{i}\rangle^{*})\langle R_{n}|b_{i}\rangle$ with $R_{n}$ and $L_{n}$ are the $n^{\text{th}}$ right and left eigenstates respectively, and $|b_{i}\rangle$’s are the Fock space chosen as computational basis. The average PR is defined as: $P^{\rm LR}_{\rm avg}=\sum_{n}P^{\rm LR}_{n}/N_{\rm L}$. We check that such a result is also consistent without the $LU$ decomposition.
Figure 2: Average of participation ratio ($P_{\text{avg}}$), using biorthonormal basis considering both left and right eigenvectors of the Liouvillian for dissipative ADM as a function of the rotating and counter-rotating coupling parameters $g_{1}$ and $g_{2}$. The atom number: $N=10$, the bosonic cut-off: $n_{\text{max}}=26$. For thisfigure $\omega_{0}=\omega=1,\ \kappa=1$.
In Fig. 2, we show the value of $P_{\rm avg}^{\rm LR}$ as a function of the coupling strengths. In the dark portion, the average participation ratio is relatively low, which indicates the non-ergodic phase, compared to the light color regime, where the average participation is quite high and represents the ergodic phase. The phase diagram is similar to Fig. 7 in the Appendix- A, where we show the transition from the non-ergodic to the ergodic phase with the aid of the level spacing ratio of eigenvalues of $\mathcal{L}$. Hence, we conclude that the eigenvector properties of our system are consistent with the eigenvalue properties. Our result also indicates that dissipation does not destroy the non-erogodicity in the system in the vicinity of $g_{1}=0$ or $g_{2}=0$. We also provide supporting plots in Appendix C to demonstrate that the qualitative features remain unchanged as the dissipation rate $\kappa$ is varied.
@@
Figure 3: (a, b, e, f, i, j, m, n) Average boson number, $N_{\text{av}}(t)$, and (c, d, g, h, k, l, o, p) mutual information between spins and bosons, $I_{\text{spin-boson}}$, as a function of time $t_{n}=2^{n}T$ for the dissipative ADM under the Thue-Morse quasiperiodic drive. The initial states have low energies, so that $\langle E_{\text{in}}\rangle=0.25$. Panels (a, c, e, g, i, k, m, o) represent the data for a fixed driving frequency $\omega_{\text{d}}=50$, $g_{1}=1.25$, and various values of $g_{2}$. Panels (b, d, f, h, j, l, n, p): $g_{1}=1.25$, $g_{2}=0.7$ and various values $\omega_{\text{d}}$. In all panels, the driving amplitude is $\Omega=1.0$, $N=6,\ n_{\text{max}}=20,\ \omega_{0}=\omega=1$. For (a-d) $\kappa=0.0$, (e-h) $\kappa=0.05$, (i-l) $\kappa=0.6$, (m-p) $\kappa=1.0$. The time t is in units of $\omega_{0}^{-1}$. In this figure, the dashed line represents the page value
IV Analysis for the driven case
In our earlier study of the closed version of driven ADM Das et al. (2023b), we show that under the external periodic drive, the system reaches a long-lived prethermal plateau that escapes Floquet heating. In contrast, for the quasi-periodic drive, the system features a short-lived prethermal plateau, and further heating is observed. At the same time, the dynamics shows the following properties,
  1. 1.
    With increasing the driving frequency, the heating process becomes slow.
  2. 2.
    In the non-ergodic regime, for the same driving frequency, heating occurs more slowly than in the ergodic regime.
Such analysis can also be extended to the dissipative case. In case of strong bosonic dissipation, the system will reach ground state where the mutual information between the spins and boson is zero. Whereas the quasiperiodic drive pushes the system to a state with higher mutual information. Hence, a competition between the quasiperiodic drive and the bosonic dissipation can be expected in this case.
IV.1 Driving protocol
@@
We briefly discuss the quasi-periodic driving protocol. In particular, we consider the quasi-periodic protocol consisting of the Thue-Morse sequence. Here, we define two Hamiltonian $H_{\pm}=\omega a^{\dagger}a+\omega_{0}J_{z}+\frac{(g_{1}\pm\Omega)}{\sqrt{N}}(aJ_{+}+a^{\dagger}J_{-})+\frac{(g_{2}\pm\Omega)}{\sqrt{N}}(aJ_{-}+a^{\dagger}J_{+})$, where $\Omega$ is the driving amplitude and the corresponding Liouvillian using Eq. (4) are $\mathcal{L}_{\pm}$. The Thue-Morse sequence Thue (1906); Nandy et al. (2017); Mukherjee et al. (2020); Zhao et al. (2021, 2022) can be constructed by starting time evolution operators $\mathcal{U}_{\pm}=\exp(\mathcal{L}_{\pm}T)$. We can build up the next sequence using $\mathcal{U}_{1}=\mathcal{U}_{-}\mathcal{U}_{+}$ and $\tilde{\mathcal{U}}_{1}=\mathcal{U}_{+}\mathcal{U}_{-}$, $\mathcal{U}_{2}=\tilde{\mathcal{U}}_{1}\mathcal{U}_{1}$ and $\tilde{\mathcal{U}}_{2}=\mathcal{U}_{1}\tilde{\mathcal{U}}_{1}$ and so on. The driving unit cells of time length $2^{n}T$ can be recursively constructed as: $\mathcal{U}_{n+1}=\tilde{\mathcal{U}}_{n}\mathcal{U}_{n},\ \tilde{\mathcal{U}}_{n+1}=\mathcal{U}_{n}\tilde{\mathcal{U}}_{n}$. The time evolution of the initial density matrix is given by $\rho_{n}|_{N_{\text{L}}\times 1}=\rho(t=2^{n}T)|_{N_{\text{L}}\times 1}=\mathcal{U}_{n}|_{N_{\text{L}}\times N_{\text{L}}}\times\rho_{\text{in}}|_{N_{\text{L}}\times 1}$, where $\rho_{\text{in}}|_{N_{\text{L}}\times 1}=|\psi_{\text{in}}\rangle|_{N_{\text{D}}\times 1}\otimes|\psi_{\text{in}}\rangle|_{N_{\text{D}}\times 1}$ is the initial density matrix and $|\psi_{\text{in}}\rangle$ is the initial state. We also define $\omega_{d}=2\pi/T$. To calculate the expectation value of an observable, we reshape the density matrix as: $\rho(t)|_{N_{\text{L}}\times 1}\rightarrow\rho(t)|_{N_{\text{D}}\times N_{\text{D}}}$. Here we take the average over $20$ low-energy eigenstates (starting from the ground state) of the decoupled Hamiltonian ($g_{1}=g_{2}=0.0$).
IV.2 Effect of the dissipation in Floquet heating
@@
To study quantum dynamics, we define the following relevant observables: the average boson number:
$N_{\text{av}}(t)=\text{Tr}\left[\rho(t)|{N{\text{D}}\times N_{\text{D}}}\times(a^{\dagger}a)|{N{\text{D}}\times N_{\text{D}}}\right]$ (9)
@@
where $S_{\text{spin-boson}}=-\text{Tr}\left[\rho(t)\ln(\rho(t))\right]$, $S_{\text{spin(boson)}}=-\text{Tr}\left[\rho_{\text{spin(boson)}}(t)\ln(\rho_{\text{spin(boson)}}(t))\right]$ are the corresponding entropies and $\rho_{\text{spin}}(t)|_{(N+1)\times(N+1)}=\text{Tr}_{\text{boson}}\left[\rho(t)|_{N_{\text{D}}\times N_{\text{D}}}\right]$, $\rho_{\text{boson}}(t)|_{(n_{\text{max}}+1)\times(n_{\text{max}}+1)}=\text{Tr}_{\text{spin}}\left[\rho(t)|_{N_{\text{D}}\times N_{\text{D}}}\right]$ is the corresponding reduced density matrix of the spin(boson) obtained by tracing over the bosonic(spin) degrees of freedom. In Fig. 3, we consider low-energy initial states subjected to the Thue–Morse driving protocol. We present the time evolution of the average boson number [Fig. 3(a, b, e, f, i, j, m, n)] and the mutual information between spins and bosons [Fig. 3(c, d, g, h, k, l, o, p)]. These quantities are shown for a fixed intermediate driving frequency $\omega_{\text{d}}$ and various values of the coupling parameter $g_{2}$ [Fig. 3(a, c, e, g, i, k, m, o)]. In addition, we analyze the same observables for a fixed $(g_{1},g_{2})$ pair corresponding to the chaotic undriven model while varying the driving frequency $\omega_{\text{d}}$ [Fig. 3(b, d, f, h, j, l, n, p)].
The key observations in this case are given below.
  1. 1.
    Fig. 3(a-d) corresponds to $\kappa=0$, and shows a result similar to the closed quasiperiodic driven ADM.
  2. 2.
    For small values of $\kappa=0.05$, we show that further heating is minimized, since both $I_{\text{spin-boson}}$, $N_{av}(t)$ are saturated less than the page value (Fig. 3(e-h)).
  3. 3.
    For increasing $\kappa$, the saturation value is further reduced. At the same time, for the non-ergodic phase, this saturation value is less than its ergodic counterpart (Fig. 3(i-p)).
  4. 4.
    For the higher $\kappa$ regime, the saturation value and the corresponding saturation time are independent of $\omega_{d}$ (Fig. 3(n,p)).
In summary, dissipation plays a constructive role in stabilizing the long-lived prethermal phase. In the next section, we also discuss the possible implications of our results for the rectification of an erroneous Floquet drive.
V Outlook
As a summary, we present a brief analysis of the existing phases in ADM in the presence of dissipation. The QPT in this case is modified by anisotropy $(g_{1}\neq g_{2})$ and finite dissipation $(\kappa>0)$. The ADM also shows ENET, as the vicinity of $g_{1}=0$, or $g_{2}=0$ shows the notion of non-ergodic phase, and the remaining part shows the ergodic behavior. In this work, we show that dissipation doesn’t destroy the existence of ENET. We support our claim with the study of the Liouvillian ($\mathcal{L}$) of the atom+cavity. The Liouvillian gap ($\Delta$) shows two different scalings of ergodic and non-ergodic phases for changing the atom numbers. Similarly, the average participation ratio of the eigenvectors explicitly shows the notion of ENET.
Finally, we consider the driven ADM, where the periodic drive leads the system to the prethermal phase. A modification of the periodic Floquet drive (e.g., quasi-periodic drive) further introduces the heating process of the prethermal state. We find that the presence of dissipation is beneficial in this case. It is known that reservoir engineering protocol can be used for the ground state or targeted state preparation, as the unwanted excitation can be released through the leaking cavity to the environment. Motivated by this fact, we aim to investigate a similar situation here. Our result indicates that dissipation halts the excessive heating for the quasi-periodic case (Fig. 3(e-p)). For the experimental realization, we need to design a driving protocol that interpolates between the two types of drives. $f(t)=a_{1}\cos(\omega_{d}t)+a_{2}\cos(\alpha\omega_{d}t)$. Here $\alpha$ is the irrational number. In the case of nuclear magnetic resonance, such a quasi-periodic drive is already implemented by providing two periodic delta kicks in the ‘y’ direction, with $\alpha=(\sqrt{5}+1)/2$ to observe discrete time quasi-crystals He et al. (2025).
We also present possible future directions for our research. An extension of our analysis to TPT and ESQPT that incorporates dissipation is required. Similarly, the study of discrete time crystals in dissipative ADM under periodic driving would be of interest for understanding dynamical phases. Furthermore, the rapidly increasing dimensionality of the Liouvillian poses a challenge for simulating large atomic ensembles; addressing this issue will require the development of advanced numerical techniques in future work.